### v1

# SOC-DGL: Social Interaction Behavior Inspired Dual Graph Learning Framework for Drug-Target Interaction Identification

Abstract
The identification of drug-target interactions (DTI) is crucial for drug discovery and repositioning, as it reveals potential uses of existing drugs, aiding in the acceleration of the drug development process and reducing associated costs. Despite the similarity information in DTI is important, most models are limited to mining direct similarity information within homogeneous graphs, overlooking the potential yet rich similarity information in heterogeneous graphs. Inspired by real-world social interaction behaviors, we propose SOC-DGL, which comprises two specialized modules: the Affinity-Driven Graph Learning (ADGL) module and the Equilibrium-Driven Graph Learning (EDGL) module. The ADGL module adopts a comprehensive social interaction strategy, leveraging an affinity-enhanced global drug-target graph to learn both global DTI and the individual similarity information of drugs and targets. In contrast, the EDGL module employs a higher-order social interaction strategy, amplifying the influence of even-hop neighbors through an even-polynomial graph filter grounded in balance theory, enabling the indirect mining of higher-order homogeneous information. This dual approach enables SOC-DGL to effectively and comprehensively capture similarity information across diverse interaction scales within the affinity matrices and drug-target association matrices, significantly enhancing the model’s generalization capability and predictive accuracy in DTI tasks. To address the issue of imbalance in drug-target interaction datasets, this paper proposes an adjustable imbalance loss function that mitigates the impact of sample imbalance by adjusting the weight of negative samples and a parameter. Extensive experiments on four benchmark datasets demonstrate significant accuracy improvements achieved by SOC-DGL, particularly in scenarios involving data imbalance and unseen drugs or targets.
{IEEEkeywords}
Drug-target interaction, affinity-driven graph learning, equilibrium-driven graph learning, data imbalance.
1 Introduction
[…]
Drug repurposing, the strategy of applying approved medications to novel therapeutic indications, has gained prominence due to the stagnation in traditional drug discovery [1]. Drug-target interactions (DTI) are central to pharmacology, yet conventional drug discovery remains time-consuming, costly, and high-risk [2]. In contrast, repurposing seeks to identify new uses for existing drugs, significantly reducing the time and cost of drug development [3]. Consequently, computational models for drug-target interactions have emerged, utilizing previous experimental data to predict novel DTI strengths. With the expanding availability of biomedical data, artificial intelligence-based methods for drug repurposing have proliferated  [4] , primarily falling into two key categories: machine learning-based methods and increasingly popular network-based methods.
Predicting DTI typically involves mapping drug-target pairs to feature vectors and applying machine learning for classification [5, 6, 7]. However, machine learning-based methods are limited by the need for positive and negative samples and are susceptible to sample imbalance, which can introduce noise and reduce prediction accuracy. Network-based methods typically build homogeneous or heterogeneous biological networks, infer interactions from drug-drug similarity, target-target similarity, or DTI, and predict interactions via network reasoning [8, 9, 10, 11]. Although network-based reasoning methods are effective to some extent, they primarily focus on direct interactions and fail to explore similarity information embedded in heterogeneous information networks.
To address this challenge, this study proposes a novel social interaction-inspired dual-enhanced graph reasoning mechanism, termed SOC-DGL, which integrates ADGL and EDGL for accurate and robust DTI prediction based on similarity information from diverse sources. The main contributions of this study are:
  • Inspired by real-world social interactions, this study proposes two complementary modules, ADGL and EDGL. Based on ADGL, which is built upon an extensive social interaction framework to achieve global representation learning on affinity-enhanced drug-target networks, EDGL further incorporates balance theory and employs an even-polynomial graph filter to capture high-order even-hop neighbor relationships in the heterogeneous drug-target graph. These two modules work in a complementary manner, enabling a comprehensive exploration of homogeneous information for both drugs and targets across multiple interaction scales.
  • To mitigate the impact of imbalanced data on the performance of SOC-DGL, this study designs a novel an adjustable imbalance loss function that mitigates the impact of sample imbalance by weighting negative samples and introducing an adjustable parameter $\varpi$. Extensive experimental results on imbalanced benchmark datasets consistently demonstrate the effectiveness.
  • ABL1, a key pathogenic target in Chronic Myeloid Leukemia, is effectively targeted by 10 drugs predicted by SOC-DGL. Utilizing a homogeneous information aggregation mechanism, SOC-DGL uncovers a potential association between Debromohymenialdisine and ABL1, highlighting its capability to identify novel DTI.
2 Related Work
2.1 DTI Prediction Based on Similarity
The integration of similarity information plays a pivotal role in advancing DTI prediction models. Traditionally, similarity mining has focused on isomorphic networks, where pairwise similarity between drugs and targets is utilized [12, 13, 14]. However, recent work has shifted towards heterogeneous graphs, which integrate multiple types of nodes and relationships [15, 16, 17]. Despite these advancements, these approaches primarily rely on similarity information derived from traditional similarity matrices, overlooking the rich and potentially valuable indirect similarity information between drugs and targets embedded within heterogeneous graphs. Such higher-order similarity information is critical for predicting potential DTI associations and necessitates the design of specialized modules for its learning.
2.2 High-order Similarity Information Learning
Effectively learning high-order similarity information aids in uncovering complex higher-order isomorphic relationships in graph data. Mi et al.[18] proposed HSL, which learns high-order similarity by considering both first-order and high-order similarities in a low-dimensional embedding space. Wang et al.[19] introduced HSDA for speech emotion recognition, capturing high-order similarity through a similarity graph to reveal cross-domain structural relationships. Peng et al. [20] proposed HSC, where high-order similarity is learned from relationships between dimensional spaces and jointly optimized with ordinary similarity for clustering. Inspired by these methods and balance theory[21], we propose the EDGL module, built upon the ADGL module, to fully exploit high-order similarity information for DTI prediction.
3 Datasets and Methodology
3.1 Datasets
To comprehensively assess the performance and applicability of SOC-DGL, this study selects four representative benchmark datasets: KIBA [22], Davis [23], BindingDB [24], and DrugBank [25], whose detailed description appears in Appendix Section A.
3.2 Model framework overview
The framework of SOC-DGL is composed of six main steps, as shown in Fig. 1, including: (I) Affinity-Enhanced Global Drug-Target Network Construction, (II) Affinity-Driven Graph Learning, (III) Equilibrium-Driven Graph Learning, (IV) Feature Fusion, (V) DTI Prediction Decoder, and (VI) Imbalanced Processing.
Figure 1: The framework of SOC-DGL. (A) Affinity-enhanced global drug-target network construction, (B) Affinity-Driven Graph Learning, (C) Equilibrium-Driven Graph Learning, (D) Feature fusion, (E) DTI prediction decoder, and (F) Imbalanced Processing.
3.2.1 Affinity-Enhanced Global Drug-Target Network Construction
In DTI prediction, diverse features of drugs and targets are derived from various measurement methods, offering complementary perspectives on molecular characteristics. Integrating these heterogeneous features is crucial for accurate DTI inference. Assuming a shared latent space for drug and target features, SOC-DGL employs pairwise multi-view learning to derive a unified subspace representation from multiple views, specifically the drug affinity matrix ($A_{DD}$) and target affinity matrix ($A_{TT}$). SOC-DGL effectively integrates multi-view information by constructing a protocol matrix that combines correlation matrices under low-rank and sparsity constraints. Details are provided in Appendix Section B.
The drug-target affinity matrix is denoted as $A_{DT}$. By combining $A_{DD}$, $A_{TT}$, and $A_{DT}$, this study constructs an affinity-enhanced global drug-target network matrix $H$, where $A_{TD}$ is the transpose of $A_{DT}$, and $H$ is a symmetric matrix:
$H=\begin{pmatrix}A_{DD}&A_{DT}\\ A_{TD}&A_{TT}\end{pmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times(n_{d}+n_{t})}$ (1)
3.2.2 Affinity-Driven Graph Learning
To extract both direct and global DTI and similarity information of drugs and targets, this study employs an extensive social interaction strategy and apply global graph convolutional learning to the affinity-enhanced global drug-target network matrix.
To effectively filter the useful information within the heterogeneous graph, for drug affinity matrix $A_{DD}$ and the target affinity matrix $A_{TT}$, elements in these two affinity matrices that are greater than or equal to the threshold 0.8 are set to 1, and the rest are set to 0, resulting in the drug affinity adjacency matrix $\tilde{A}_{DD}$ and the target affinity adjacency matrix $\tilde{A}_{TT}$. By combining the drug-target affinity matrix, this study constructs the adjacency matrix $G$ of:
$G=\begin{pmatrix}\tilde{A}_{DD}&A_{DT}\\ {A}_{TD}&\tilde{A}_{TT}\end{pmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times(n_{d}+n_{t})}$ (2)
Subsequently, this study performs graph convolution operations:
$H^{(l+1)}=\sigma(D^{-\frac{1}{2}}GD^{-\frac{1}{2}}H^{(l)}W^{(l)})$ (3)
where $D^{-1/2}GD^{-1/2}$ represents the normalization process, with $D=\mathrm{diag}\left(\sum_{j}G_{ij}\right)$ being the degree matrix of the adjacency matrix $G$. $H^{(l)}$ is the feature embedding matrix at layer $l$, $W^{(l)}$ is the learnable weight matrix at layer $l$, and $\sigma$ is the activation function. After two rounds of graph convolution, the graph convolution feature embedding matrix $H^{\prime}$ is obtained:
$H^{\prime}=\begin{bmatrix}H^{\prime}_{D}\\ H^{\prime}_{T}\end{bmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times m}$ (4)
where $H^{\prime}_{D}\in\mathbb{R}^{n_{d}\times m}$ is the drug feature embedding matrix, $H^{\prime}_{T}\in\mathbb{R}^{n_{t}\times m}$ is the target feature embedding matrix, and $m$ is the final output dimension.
3.2.3 Equilibrium-Driven Graph Learning
In the real situation, drugs with similar structures may exhibit analogous mechanisms of action, while the similarity between targets reflects the functional conservation of targets [26]. As the classic adage goes: “The friend of my friend is also my friend,” this concept is equally applicable to DTI. Using balance theory [21], this study deduces that even though drugs and targets differ in their attributes, they can establish connections through shared interaction relationships. If a drug interacts with multiple targets, and these targets are connected to other drugs, then these drugs and targets may form a state of equilibrium through these indirect interaction relationships. To this end, this study proposes the EDGL module, which employs a higher-order social interaction strategy to indirectly learn higher-order drug and target similarity information linked by ’friend nodes’ through an even-polynomial graph filter.
This study processes $H^{\prime}$, which has been encoded through ADGL. For ease of expression, let $X_{0}=H^{\prime}$, and then perform two linear transformations on $X_{0}$, applying an activation function. $X$ is obtained as follows:
$X_{(l+1)}=\sigma(X_{(l)}\cdot W_{(l)}+b_{(l)})$ (5)
This study normalizes the adjacency matrix $G$ by setting $G^{\prime}=D^{-1/2}GD^{-1/2}$, where $D$ is the degree matrix of the adjacency matrix $G$. Thus,
$G^{\prime}=\begin{pmatrix}\tilde{A}_{DD}^{\prime}&A_{DT}^{\prime}\\ A_{TD}^{\prime}&\tilde{A}_{TT}^{\prime}\end{pmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times(n_{d}+n_{t})}$ (6)
To learn similarity information in drug-target association matrix, this study sets $(\tilde{A}_{DD}^{\prime})_{ij}=0$, $(\tilde{A}_{TT}^{\prime})_{ij}=0$, thereby transforming it into the adjacency matrix $A^{\prime}$:
$A^{\prime}=\begin{pmatrix}0&A_{DT}^{\prime}\\ A_{TD}^{\prime}&0\end{pmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times(n_{d}+n_{t})}$ (7)
where $A_{DP}^{\prime}\in\mathbb{R}^{n_{d}\times n_{t}}$, $\tilde{A}_{DD}^{\prime}\in\mathbb{R}^{n_{d}\times n_{d}}$. This study then obtains the Laplacian matrix $L=D^{\prime}-A^{\prime}$, where $D^{\prime}$ is the degree matrix of $L$. After standardizing $L$, the formula is as follows:
$\tilde{L}=D^{\prime-\frac{1}{2}}LD^{\prime-\frac{1}{2}}=I-D^{\prime-\frac{1}{2}}A^{\prime}D^{\prime-\frac{1}{2}}$ (8)
where $I$ is the identity matrix. Subsequently, this study defines the propagation matrix as $P=I-\tilde{L}=D^{-1/2}A^{\prime}D^{-1/2}$, and the $k$-th order polynomial graph filter is defined as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}\tilde{L}^{k}$, $k=0,1,\ldots,K$, where $\alpha_{k}$ is the learnable weight matrix, and $\alpha_{k}=\alpha\times(1-\alpha)^{k}$, with $\alpha$ being a hyperparameter that can be set. Thereafter, this study rewrites the graph filter formula as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}(I-\tilde{L})^{k}=\sum_{k=0}^{K}\alpha_{k}P^{k}$, discarding the odd-degree terms in $g(\tilde{L})$ and retaining only even-degree terms:
$g_{\text{even}}(\tilde{L})=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{2k}(I-\tilde{L})^{2k}=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k}$ (9)
Thus, the model of EDGL is constructed:
$Y=f(\hat{X},P)=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k}\hat{X}$ (10)
where $\hat{X}$ is the input feature matrix that has already undergone two linear transformations, and $Y$ represents the final output features. This study denotes the result obtained after passing $H^{\prime}$ through EDGL as $H^{\prime\prime}$, and its expression is as follows:
$H^{\prime\prime}=Y=\begin{bmatrix}H_{D}^{\prime\prime}\\ H_{T}^{\prime\prime}\end{bmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times m}$ (11)
3.2.4 Feature Fusion
This study employs residual connection techniques to integrate information from diverse sources, ensuring that the two modules complement each other and collectively form a more powerful graph reasoning mechanism. ”This approach balances the influence of higher-order interaction patterns on the generalized patterns, achieving equilibrium between the two ”social interaction paradigms.” The corresponding formula is as follows:
$\hat{H}=\omega H^{\prime}+(1-\omega)H^{\prime\prime}$ (12)
where $\omega$ represents the weight, which is an adjustable hyperparameter. The integrated features $\hat{H}$ will serve as the input for the decoder, used to predict DTI. Through residual connection feature fusion, this study gains a multifaceted understanding of the complex relationships between drugs and targets.
3.2.5 DTI prediction decoder
DTI prediction decoder in SOC-DGL cleverly employs the principle of matrix factorization. By meticulously decoding these features, it reveals the affinity between drugs and targets, assessing the potential interaction affinity between different drugs and targets. This study denotes $W^{L}$ as the learnable weight matrix for the mapping from the drug space to the target space within the decoder. Inspired by matrix tri-factorization, the formula for DTI prediction decoder is as follows:
$H^{*}=\sigma(\hat{H}_{D}W^{(L)}\hat{H}_{P}^{T})$ (13)
where $H^{*}$ represents the final predicted interaction matrix, and its element $H_{ij}^{*}$ denotes the predicted association score between drug $i$ and target $j$.
3.2.6 Imbalanced Processing
To address the issue of class imbalance between positive and negative samples in the model, this study designs a Weighted Binary Cross-Entropy Loss Function with Reduced Regative Samples (RLF), as shown in the following formula:
$\mathcal{L}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (14)
where $n_{d}$ is the number of drugs, $n_{t}$ is the number of targets, $|y^{+}|$ represents the number of positive samples, and $|y^{-}|$ represents the number of negative samples. $h_{i}^{*}$ represents the model’s predicted probability of interaction between drug $i$ and target $j$. This novel loss function is built upon the existing weighted cross-entropy loss function, with the introduction of an adjustable parameter $\varpi$, which effectively reduces the weights of negative samples while keeping the weights of positive samples unchanged.
Table 1: SOC-DGL performance in the “warm-start” scenario on four benchmark datasets.
Dataset KIBA dataset Davis dataset BindingDB dataset DrugBank dataset
Situation balanced imbalanced balanced imbalanced balanced imbalanced balanced imbalanced
AUROC 0.9761 $\pm$ 0.0023 0.9364 $\pm$ 0.0036 0.9467 $\pm$ 0.0074 0.9192 $\pm$ 0.0045 0.9894 $\pm$ 0.0032 0.9790 $\pm$ 0.0034 0.9735 $\pm$ 0.0023 0.9295 $\pm$ 0.0076
AUPR 0.9805 $\pm$ 0.0018 0.7473 $\pm$ 0.0123 0.9673 $\pm$ 0.0048 0.7655 $\pm$ 0.0134 0.9888 $\pm$ 0.0036 0.8705 $\pm$ 0.0146 0.9774 $\pm$ 0.0019 0.7654 $\pm$ 0.0126
F1_score 0.9350 $\pm$ 0.0041 0.6908 $\pm$ 0.0068 0.9038 $\pm$ 0.0103 0.6918 $\pm$ 0.0129 0.9641 $\pm$ 0.0075 0.8057 $\pm$ 0.0177 0.9286 $\pm$ 0.0027 0.7211 $\pm$ 0.0104
ACC 0.9268 $\pm$ 0.0049 0.9207 $\pm$ 0.0026 0.8816 $\pm$ 0.0131 0.9134 $\pm$ 0.0030 0.9627 $\pm$ 0.0077 0.9594 $\pm$ 0.0041 0.9267 $\pm$ 0.0026 0.9508 $\pm$ 0.0015
Recall 0.9373 $\pm$ 0.0093 0.7103 $\pm$ 0.0182 0.9097 $\pm$ 0.0200 0.6642 $\pm$ 0.0338 0.9677 $\pm$ 0.0098 0.8406 $\pm$ 0.0246 0.9208 $\pm$ 0.0136 0.6567 $\pm$ 0.0210
Specificity 0.9133 $\pm$ 0.0148 0.9506 $\pm$ 0.0049 0.8372 $\pm$ 0.0362 0.9561 $\pm$ 0.0066 0.9574 $\pm$ 0.0107 0.9726 $\pm$ 0.0047 0.9331 $\pm$ 0.0130 0.9824 $\pm$ 0.0022
Precision 0.9328 $\pm$ 0.0100 0.6730 $\pm$ 0.0177 0.8985 $\pm$ 0.0189 0.7238 $\pm$ 0.0230 0.9606 $\pm$ 0.0097 0.7744 $\pm$ 0.0304 0.9369 $\pm$ 0.0109 0.8004 $\pm$ 0.0174
Figure 2: Performance comparison of SOC-DGL with baseline methods on the KIBA dataset. (A) ROC and PR curves on balanced datasets and (B) imbalanced datasets. (C) F1 score, ACC, recall, specificity, and precision on balanced datasets and (D) imbalanced datasets.
4 Experiments and Results
4.1 Experimental design overview
In the ”warm-start” scenario, this study designed validation experiments for both balanced datasets (positive:negative samples = 1:1) and imbalanced datasets (positive:negative samples = 1:10). In most real-world tasks, the ratio of positive to negative samples is typically imbalanced. In drug-target association prediction, known associations (positive samples) were generally far fewer than unknown associations (negative samples). Evaluating with imbalanced datasets better reflected SOC-DGL’s performance in real-world settings. Conversely, balancing the positive and negative samples helped mitigate model bias towards predicting negative samples, providing a more balanced assessment of model performance.
To minimize the impact of data variability on the results, this study employed 10-fold cross-validation to assess the model’s predictive performance. First, the entire dataset (positive and negative samples) was randomly divided into 10 subsets of approximately equal size. This process was repeated 10 times, with each time selecting a different subset as the test set and the other 9 subsets as the training set. Initially, the test set data was masked, and then the model was trained on the remaining 9 subsets. The masked test set values were predicted for DTI, and the predicted values were compared with the actual DTI values (before masking) using the AUROC, AUPR, F1_score, Accuracy (ACC), Recall, Specificity, and Precision metrics.
In this study, in the “warm-start” scenario, comprehensive experimental validation was conducted on four representative benchmark datasets to assess the performance of SOC-DGL. For other specific experiments, the study primarily chose to conduct validation on the KIBA dataset due to its moderate data size and comprehensive information, which provided accurate support for model training and evaluation.
This study conducted experiments using PyTorch 1.11.0 on an Nvidia GeForce RTX 3060 GPU, and employed the Xavier initialization strategy to prevent gradient explosion and vanishing during the training process.
4.2 “Warm-start” experiments
In this study, “Warm-start” experiments were conducted on four benchmark datasets under both balanced and imbalanced settings. Multiple runs were performed on each dataset with optimized parameter configurations, and average metrics with standard deviations were reported, as shown in Tab. 1. The results demonstrated that SOC-DGL achieved high predictive accuracy in balanced settings, effectively distinguishing known DTIs, and performed well in the “warm-start” scenario, validating application in DTI prediction and drug discovery.
To assess SOC-DGL’s effectiveness, comparative experiments were conducted with four leading DTI prediction methods: GraphDTA [27], LRSpNM [28], MLMC [29], MULGA [30], and MMDG-DTI [31], representing the state-of-the-art. Detailed comparisons were available in Appendix Section C. SOC-DGL showed significant improvements on both balanced and imbalanced datasets, demonstrating its broad applicability and outperformance over current methods.
As shown in Fig. 2(A) and Fig. 2(B), SOC-DGL consistently outperformed all baseline methods on both balanced and imbalanced datasets. It was clearly evident that MMDG-DTI consistently performed just slightly behind SOC-DGL on the KIBA balanced dataset. MLMC also showed strong performance, leveraging multi-view learning and matrix completion techniques. In contrast, LRSpNM, based on low-rank matrix approximation, performed the worst across all datasets, likely due to DPI matrix sparsity and inaccurate similarity measures. On the KIBA balanced dataset, SOC-DGL’s AUROC and AUPR were 1.8% and 1.6% higher compared to the current best-performing model, respectively, and 0.4% and 0.1% higher in the imbalanced setting. MLMC and MULGA performed slightly worse than SOC-DGL, with MULGA underperforming in imbalanced settings due to inadequate attention to homogeneous association information. However, in other datasets, MULGA’s multi-view learning approach combined with graph autoencoders ranked second to SOC-DGL, suggesting lower generalizability. Appendix Tab. I further confirmed SOC-DGL’s superiority across datasets. The F1-score, ACC, Recall, Specificity, and Precision results in Fig. 2(C) and Fig. 2(D) further validated SOC-DGL’s consistent outperformance.
Figure 3: Performance scores of SOC-DGL in ”cold-start” scenarios for drugs and targets, evaluated on various metrics. Black dots represent randomly sampled data points corresponding to indicator values of these metrics.
4.3 “Cold-start” experiments
In the “cold-start” experiments on the KIBA dataset, two scenarios were designed: “cold-start for drugs” and “cold-start for targets,” where the drug-target affinity matrix was systematically sampled. In the “cold-start for targets” scenario, drugs linked to specific targets were removed, treating each target as new, with the removed data serving as the test set. In the “cold-start for drugs” scenario, drugs linked to specific targets were removed, and the corresponding test set consisted of the removed drugs and their targets.
The KIBA dataset, comprising 1,720 drugs and 220 targets, provided the foundation for “Cold-start” experiments. A total of 56 independent cold-start experiments were conducted, involving the random selection of 34 drugs and 22 targets to simulate real-world scenarios where new drugs or targets lacked historical interaction data. This was particularly relevant in drug discovery, where novel compounds or under-explored proteins often lacked prior interaction data.
Fig. 3 presented the results. In the “cold-start for drugs” scenario, SOC-DGL showed stable performance across different drugs, with an average AUROC of 0.9758 and AUPR of 0.9803, indicating strong generalization ability for new drugs. Similarly, in the “cold-start for targets” scenario, SOC-DGL demonstrated consistent performance across different targets. These results confirmed its robustness in predicting drug-target interactions for unseen drugs or targets, making it a valuable tool for practical drug development and target discovery.
4.4 Analysis of the Even-Polynomial Graph Filter in EDGL
To analyze the importance of the even-polynomial graph filter in EDGL for indirectly extracting higher-order homogeneous information between drugs and targets, this study compared it with the odd-polynomial graph filter for indirectly extracting higher-order heterogeneous information and with the self-attention mechanism under the “warm start” scenario, while keeping all other modules unchanged. Specifically, this study evaluated the effectiveness of higher-order similarity information enhancement by the even-polynomial graph filter (Homogeneous_prop), higher-order interaction information enhancement by the odd-polynomial graph filter (Heterogeneous_prop), and self-information enhancement by the self-attention mechanism (Self_attention_prop) using various metrics to comprehensively measure their performance.
As shown in Fig. 4(A), Homogeneous_prop outperformed Heterogeneous_prop and Self_attention_prop, particularly in AUROC and AUPR, excelling in DTI prediction and handling imbalanced data by capturing complex indirect relationships. While Heterogeneous_prop was effective for direct connections, it struggled with indirect ones. Self_attention_prop, despite its node-weighting capability, performed weaker overall in managing complex interactions. Homogeneous_prop showed a trade-off with slightly lower Recall, indicating reduced sensitivity to some positives. These results highlighted the superior ability of Homogeneous_prop to capture indirect interactions and improve prediction accuracy.
4.5 Ablation Study on Diversity Graph Learning
To evaluate the impact of different graph learning modules, three SOC-DGL variants were tested: (1) ADGL-Only, focusing on global associations, (2) EDGL-Only, emphasizing higher-order similarities, and (3) Fusion-Without, combining the two modules without feature fusion residual connection to assess the role of feature fusion.
As shown in Fig. 4 (B), SOC-DGL outperformed the variants on most metrics, particularly in AUROC and AUPR. On balanced datasets, SOC-DGL achieved an AUROC of 0.9761, surpassing ADGL-Only (0.9660) and EDGL-Only (0.9703), demonstrating the clear benefit of combining ADGL and EDGL for capturing complex DTIs. Similarly, SOC-DGL’s AUPR of 0.9805 exceeded ADGL-Only (0.9707) and EDGL-Only (0.9690). While Fusion-Without showed comparable performance in some cases, it fell short overall, indicating that the residual fusion method, which effectively balanced the importance of different DTI patterns, more effectively identified and integrated complex interaction pathways, thereby improving model generalization.
Figure 4: Experimental validation conducted on the KIBA dataset. (A) Experimental evaluation of Homogeneous_prop, Heterogeneous_prop, and Self_attention_prop enhancement components in SOC-DGL for DTI prediction. (B) Ablation study of different components within the SOC-DGL model.
Table 2: Warm start validation experiment on four benchmark imbalanced datasets using different loss functions.
Dataset Loss AUROC AUPR F1_score ACC Recall Specificity Precision
KIBA dataset SLF 0.9329 $\pm$ 0.0042 0.7300 $\pm$ 0.0111 0.6758 $\pm$ 0.0090 0.9156 $\pm$ 0.0036 0.7054 $\pm$ 0.0265 0.9455 $\pm$ 0.0071 0.6501 $\pm$ 0.0230
FLF 0.8983 $\pm$ 0.0078 0.6209 $\pm$ 0.0029 0.5400 $\pm$ 0.0023 0.8653 $\pm$ 0.0145 0.6530 $\pm$ 0.0175 0.8130 $\pm$ 0.0187 0.5411 $\pm$ 0.0016
WLF 0.9309 $\pm$ 0.0035 0.7081 $\pm$ 0.0074 0.6675 $\pm$ 0.0056 0.9126 $\pm$ 0.0020 0.7032 $\pm$ 0.0263 0.9425 $\pm$ 0.0058 0.6363 $\pm$ 0.0164
RLF 0.9364 $\pm$ 0.0036 0.7473 $\pm$ 0.0123 0.6908 $\pm$ 0.0068 0.9207 $\pm$ 0.0026 0.7103 $\pm$ 0.0182 0.9506 $\pm$ 0.0049 0.6730 $\pm$ 0.0177
Davis dataset SLF 0.9171 $\pm$ 0.0046 0.7601 $\pm$ 0.0089 0.6907 $\pm$ 0.0144 0.9135 $\pm$ 0.0061 0.6618 $\pm$ 0.0234 0.9558 $\pm$ 0.0085 0.7233 $\pm$ 0.0334
FLF 0.8351 $\pm$ 0.0106 0.6934 $\pm$ 0.0094 0.6132 $\pm$ 0.0122 0.8345 $\pm$ 0.0199 0.5938 $\pm$ 0.0443 0.8687 $\pm$ 0.0292 0.6376 $\pm$ 0.0161
WLF 0.9158 $\pm$ 0.0031 0.7506 $\pm$ 0.0085 0.6790 $\pm$ 0.0038 0.9097 $\pm$ 0.0042 0.6516 $\pm$ 0.0221 0.9540 $\pm$ 0.0082 0.7102 $\pm$ 0.0592
RLF 0.9192 $\pm$ 0.0045 0.7655 $\pm$ 0.0134 0.6918 $\pm$ 0.0129 0.9134 $\pm$ 0.0030 0.6642 $\pm$ 0.0338 0.9561 $\pm$ 0.0066 0.7238 $\pm$ 0.0230
BindingDB dataset SLF 0.9684 $\pm$ 0.0064 0.8031 $\pm$ 0.0302 0.7553 $\pm$ 0.0350 0.9467 $\pm$ 0.0101 0.8162 $\pm$ 0.0311 0.9612 $\pm$ 0.0116 0.7060 $\pm$ 0.0036
FLF 0.9067 $\pm$ 0.0151 0.7535 $\pm$ 0.0167 0.7807 $\pm$ 0.0151 0.8996 $\pm$ 0.0109 0.8139 $\pm$ 0.0140 0.9003 $\pm$ 0.0128 0.6994 $\pm$ 0.0268
WLF 0.9578 $\pm$ 0.0074 0.7800 $\pm$ 0.0282 0.6899 $\pm$ 0.0405 0.9336 $\pm$ 0.0116 0.7326 $\pm$ 0.0392 0.9560 $\pm$ 0.0134 0.6564 $\pm$ 0.0685
RLF 0.9790 $\pm$ 0.0034 0.8705 $\pm$ 0.0146 0.8057 $\pm$ 0.0177 0.9594 $\pm$ 0.0041 0.8406 $\pm$ 0.0246 0.9726 $\pm$ 0.0047 0.7744 $\pm$ 0.0304
DrugBank dataset SLF 0.9150 $\pm$ 0.0094 0.7316 $\pm$ 0.0256 0.6998 $\pm$ 0.0205 0.9486 $\pm$ 0.0034 0.6158 $\pm$ 0.0267 0.9845 $\pm$ 0.0022 0.8104 $\pm$ 0.0206
FLF 0.8206 $\pm$ 0.0096 0.4512 $\pm$ 0.0077 0.5127 $\pm$ 0.0133 0.8756 $\pm$ 0.0321 0.4733 $\pm$ 0.0331 0.9511 $\pm$ 0.0388 0.5755 $\pm$ 0.0414
WLF 0.9208 $\pm$ 0.0073 0.7321 $\pm$ 0.0204 0.6795 $\pm$ 0.0164 0.9388 $\pm$ 0.0029 0.6696 $\pm$ 0.0242 0.8677 $\pm$ 0.0030 0.6897 $\pm$ 0.0249
RLF 0.9295 $\pm$ 0.0076 0.7654 $\pm$ 0.0126 0.7211 $\pm$ 0.0104 0.9508 $\pm$ 0.0015 0.6567 $\pm$ 0.0210 0.9824 $\pm$ 0.0022 0.8004 $\pm$ 0.0174
4.6 Validation Experiment on Imbalanced Loss Function
In DTI prediction, there was often an imbalance between positive and negative samples, which could affect model performance, causing it to favor the majority class while neglecting the minority. This experiment evaluated the impact of various loss functions on model performance, focusing on those designed to address class imbalance.
Four loss functions were assessed: Standard Binary Cross-Entropy Loss (SLF), Focal Loss with Dynamic Scaling (FLF), Weighted Binary Cross-Entropy Loss (WLF), and Weighted Binary Cross-Entropy Loss with Reduced Negative Samples (RLF). Experiments were conducted in a “warm-start” scenario with imbalanced datasets, evaluating DTI prediction performance across multiple metrics. Details of these loss functions could be found in Appendix Section E.
As shown in Tab. 2, the RLF loss function consistently outperformed the others across four benchmark imbalanced datasets, demonstrating its effectiveness for imbalanced DTI prediction. Although RLF did not always yield the highest scores on individual metrics, it showed stable and robust performance. By incorporating a weighting parameter $\varpi$ to reduce the weight of negative samples while maintaining the positive sample weight, RLF addressed class imbalance while preserving the model’s discriminative ability. These results, observed across four distinct imbalanced datasets, further validated RLF’s strong generalization capability.
Figure 5: SOC-DGL’s performance across AUROC, AUPR, and F1_score with different GCN layer settings on (A) four balanced datasets and (B) four imbalanced datasets. AUROC, AUPR, and F1_score of SOC-DGL under different GCN layer settings on (C) balanced datasets and (D) imbalanced datasets. (E) Hyperparameter tuning of $K$ and $\alpha$ on the KIBA balanced dataset. (E) Hyperparameter tuning of $\varpi$ on the KIBA imbalanced dataset.
4.7 Hyperparameter Tuning Experiments
To assess the model’s robustness across multiple benchmark datasets, this study focused on the hyperparameters within the ADGL and EDGL modules and performed ”warm-start” experiments to systematically optimize these parameters.
4.7.1 Optimization of ADGL Hyperparameters
The number of convolutional layers ($N$) in the ADGL module was a critical hyperparameter influencing the model’s ability to capture the topological structure of drugs and targets. To assess the impact of different layer configurations on DTI prediction, experiments were conducted on both balanced and imbalanced versions of four benchmark datasets, with performance evaluated across multiple metrics, as shown in Fig. 5 (A-D). The results indicated that a two-layer GCN consistently achieved the best performance, particularly on the BindingDB, DrugBank, and Davis datasets. For $N=2$, the model excelled in AUROC, AUPR, and F1 score. However, on the KIBA dataset, $N=1$ slightly outperformed $N=2$ in AUROC (0.9838 vs. 0.9761) and AUPR (0.9872 vs. 0.9805), suggesting that a shallower GCN might have better captured global associations in this specific dataset. Overall, the two-layer setting expanded the receptive field, enhancing message passing and model expressiveness, while higher layer counts introduced complexity, increasing the risk of overfitting.
4.7.2 Optimization of EDGL Hyperparameters
This study focused on tuning key hyperparameters within the EDGL module. Specifically, this study investigated the total number of iterations ($K$) and the influence parameter ($\omega_{\alpha}$, denoted as $\alpha$) using the KIBA dataset. To thoroughly assess the impact of these hyperparameters, this study experimented with a range of values for $K$ and $\alpha$, evaluating their effects on the performance of SOC-DGL. This study explored $K$ values between 100 and 500 and adjusted $\alpha$ within the range of 0.05 to 0.25. The results presented in Fig. 5(E) showed that optimal performance was achieved when $K=200$ and $\alpha=0.20$, yielding an AUROC of 0.9761. This showed that a balanced and optimal choice of $K$ and $\alpha$ improved performance, mitigating the negative impact of extreme values.
4.7.3 Optimization of Adjustable Parameter $\varpi$
Traditional imbalanced loss functions addressed the negative sample imbalance by adjusting the influence of positive samples based on the ratio of negative to positive samples. However, this approach did not provide an optimal solution for handling the imbalance between positive and negative samples. This seemingly reasonable correction may have resulted in an excessively large influence of positive sample loss on the total loss, leading to artificial positive sample imbalance. Therefore, an adjustable parameter was required to adjust the correction for positive samples, ensuring that the imbalance function effectively handled negative sample imbalance. This study investigated the effect of the adjustable parameter $\varpi$ on imbalance processing through a hyperparameter experiment with $\varpi\in\{0.1,0.2,0.3,0.4,0.5,0.6\}$. Results in Fig. 5 (F) showed that $\varpi=0.2$ achieved the best performance, highlighting the necessity of introducing an adjustable parameter of appropriate size to adjust the correction for positive samples.
Figure 6: Case study analysis. (A) SOC-DGL predicts the top 9 drugs targeting ABL1 and demonstrates the associations among them. (B) Verification of ADGL-Edge and EDGL-Edge between Debromohymenialdisine and ABL1. (C) Perform a molecular docking experiment between Debromohymenialdisine and ABL1. The yellow dashed lines represent hydrogen bonds.
Table 3: SOC-DGL Prediction of the Top 10 Drugs Targeting Tyrosine-Protein Kinase ABL1.
No. UniProtKB Entry Name Evidence ADGL-Edge EDGL-Edge
1 DB12010 Fostamatinib Rolf et al. [32] 4 6
2 DB00171 ATP Inming et al. [33] 2 1
3 DB03878 N-[4-Methyl-3-[4-(3-Pyridinyl)-2-Pyrimidinyl]Amino]Phenyl]-3-Pyridinecarboxamide Berman et al. [34] 1 0
4 DB08896 Regorafenib DrugBank 1 0
5 DB12267 Brigatinib Markham [35] 1 0
6 DB14989 Umbralisib DrugBank 1 0
7 DB00619 Imatinib Haberler et al. [36] 1 0
8 DB01254 Dasatinib Piccaluga et al. [37] 2 1
9 DB04868 Nilotinib DrugBank 1 0
10 DB04367 Debromohymenialdisine None 3 1
4.8 Drugs prediction for the key target ABL1 in Chronic Myeloid Leukemi
To assess the potential of SOC-DGL in discovering DTI in real-world applications, this study conducted a case study on Chronic Myeloid Leukemia (CML), a myeloproliferative neoplasm with an incidence rate of approximately 1-2 cases per 100,000 adults, accounting for around 15% of newly diagnosed adult leukemia cases [38]. The primary pathogenesis of CML is driven by the fusion of the ABL1 gene on chromosome 9, a tyrosine kinase, with the Breakpoint Cluster Region (BCR) gene on chromosome 22, resulting in the formation of a malignant target that initiates and progresses the disease [39]. The ABL1 gene encodes a tyrosine kinase enzyme that activates downstream signaling pathways, regulating cell proliferation and differentiation [40]. CML onset is closely associated with ABL1 mutations, which lead to excessive cell proliferation, making it a key pathogenic factor in the disease [41].
[…]
Notably, SOC-DGL predicted Debromohymenialdisine as the 10th-ranked drug, which is a cyclin-dependent kinase inhibitor with pharmacological activity in the treatment of osteoarthritis and Alzheimer’s disease [42]. Upon searching databases and relevant literature, this study found that the direct association between Debromohymenialdisine and ABL1 had not been confirmed by any database or publication. However, SOC-DGL’s prediction suggested a potential link between the two. As shown in Fig. 6 (B) and Tab. 3, further investigation revealed that SOC-DGL first learned three direct similarity relations between DB04367, DB12267, and DB03878 through the ADGL module, as well as an interaction between DB04367 and Q07912. Subsequently, through the EDGL module, a friendship pathway was constructed using ABL1 as an intermediary, indirectly enhancing the higher-order similarity information between DB12267 and DB03878, and thereby uncovering a potential connection between ABL1 and DB04367. To further verify the binding degree of DB04367 and ABL1 at the molecular level, we conducted molecular docking experiments on them with Pymol and AutoDockTools, shown in Fig. 6 (C). The docking results indicated that DB04367 formed intermolecular forces with the amino acid residues in ABL1, further confirming the interaction between DB04367 and ABL1. These findings highlighted the remarkable ability of SOC-DGL to identify potential DTI through two distinct drug-target friendship strategies and validated the predictions of SOC-DGL, demonstrating its reliability in recognizing potential DTI and providing valuable insights for biological research.
5 Conclusion
In this study, SOC-DGL, a novel graph learning framework was proposed for DTI prediction. Drawing inspiration from real-world social interactions, this paper designed ADGL, which captures the direct global association and similarity information between drugs and targets based on extensive social connections. Additionally, EDGL leverages an even-polynomial graph filter driven by balance theory to extract higher-order similarity information from the heterogeneous drug-target graph, akin to high-order social interactions. This dual approach enhances prediction accuracy and demonstrates robust performance under data imbalance conditions. Experimental results on four benchmark datasets demonstrate that SOC-DGL outperforms existing state-of-the-art methods in both balanced and imbalanced data settings. The success of SOC-DGL in potential DTI (between DB01254 and ABL1) prediction highlights the importance of leveraging homogenous information aggregation in capturing complex indirect relationships within biological networks. Although SOC-DGL performs well in DTI prediction, it’s performance still limited by data quality, and future work could involve incorporating more biological knowledge or employing advanced data augmentation techniques to reduce noise and bias in datasets.
Supplementary Materials
6 Appendix
6.1 Datasets
To comprehensively evaluate the performance and applicability of the proposed method, we select four representative benchmark datasets: the KIBA dataset, the Davis dataset, the BindingDB dataset, and the DrugBank dataset.
6.1.1 KIBA dataset
In the KIBA dataset, some drugs and proteins do not have direct affinity metrics. However, it still contains comprehensive bioactivity measurement data between 2,111 kinase inhibitors and 229 kinases, including but not limited to key parameters such as K(d), K(i), and IC50. Notably, DTI are considered to have binding affinity when the K(d) value is below 30 units. To further optimize and utilize this dataset, we employ the Therapeutics Data Commons (TDC), a drug discovery application tool developed by Huang et al. [43] in 2022. Through the download and filtering capabilities of the TDC tool, we successfully obtain a refined dataset covering 1,720 drugs and 220 targets. This binary processed DTI dataset comprises 22,154 DTI, providing rich information and a solid foundation for our research and model training.
6.1.2 Davis dataset
The Davis dataset is a highly valuable resource in the field of DTI research, meticulously documenting the K(d) values obtained from wet laboratory tests between 72 kinase inhibitors and 442 kinases. This dataset covers over 80% of the human catalytic protein kinase family, providing a wealth of information for understanding the bioactivity of kinase inhibitors. To further enhance the practicality of the data and the accuracy of analysis, we employ the Therapeutics Data Commons tool to process the Davis dataset. Through TDC’s advanced filtering and analysis capabilities, we successfully refine 7,320 positive samples, which encompass 68 drugs and 379 targets. This binary processed DTI dataset offers a more precise and comprehensive collection of positive samples for bioinformatics research, improving the quality of the data.
6.1.3 BindingDB dataset
BindingDB is a public dataset specifically designed to store experimentally measured binding affinities between small molecules (ligands) and their corresponding target biomolecules (targets). With an extensive dataset encompassing K(d) values for 10,665 drugs and 1,413 targets, BindingDB provides researchers with a valuable platform for gaining insights into DTI. To further optimize this data for specific research needs and enhance its usability, we employ the TDC tool and retain 9,166 DTI between 3,400 drugs and 886 targets.
6.1.4 DrugBank dataset
In ensuring the accuracy and reliability of the dataset, we implement a series of stringent quality control measures. Specifically, we conduct a thorough review of the drug molecules in the dataset, manually filtering out those drug sequences that cannot be effectively recognized through the Simplified Molecular Input Line-Entry System (SMILES) format using RDKit [44]. To address the sparsity issue of the DTI dataset, we further optimize the dataset by retaining only those drugs that interact with at least two targets. After these carefully designed selection processes, we ultimately obtain a dataset comprising 6,871 verified DTI, covering 1,821 drugs and 1,447 targets.
6.1.5 Preprocessing of datasets
For each benchmark dataset, we extract various features of drugs and targets based on the SMILES sequences of drugs and the UniProt IDs of targets. For drug sequences, we use the RDKit software package to extract Molecular Access System (MACCS) keys fingerprints [45], topological (TOP) fingerprints [46], and Morgan fingerprints [47], which represent three distinct aspects of the molecular features. For target sequences, to provide a more systematic description of target sequences and their physicochemical properties, we employ the iLearn software package to calculate amino acid composition (AAC), Moran autocorrelation (MORAN) [48], Composition/Transition/Distribution (CTD) [49], and pseudo-amino acid composition (PAAC) [50], which encompass four different aspects of target features.
In drug discovery and bioinformatics, selecting negative samples is crucial. Positive samples are drugs known to interact with targets, while negative samples are those with unknown interactions. To improve model performance, we employ an association-based negative sampling method that selects more reliable negative samples by leveraging drug similarity. If a drug’s neighbors (high-similarity drugs) interact with a target, it is likely to do so as well. Conversely, drugs with low similarity to known positives are less likely to share common targets and can be considered negative samples, reducing the model’s bias toward positive interactions.
6.2 Detailed Process of Multi-view Learning
Taking the application of multi-view learning in drug affinity matrices as an example, the specific process of multi-view learning can be detailed as follows. First, we respectively represent the MACC feature matrix, the Morgan feature matrix and the TOP feature matrix as $Y_{1}\in\mathbb{R}^{v_{1}\times n}$, $Y_{2}\in\mathbb{R}^{v_{2}\times n}$, and $Y_{3}\in\mathbb{R}^{v_{3}\times n}$, where $v_{1}$ represents the dimensionality of the MACC drug features, $v_{2}$ represents the dimensionality of the Morgan drug features, $v_{3}$ represents the dimensionality of the TOP drug features, and $n$ represents the number of drugs. Consequently, the combined drug feature matrix $Y$ can be obtained, expressed as follows:
$Y=\begin{bmatrix}Y_{1}^{T}\\ Y_{2}^{T}\\ Y_{3}^{T}\end{bmatrix}\in\mathbb{R}^{(v_{1}+v_{2}+v_{3})\times n}$ (15)
Then, we define $v=v_{1}+v_{2}+v_{3}$ to represent the total number of perspectives. Following this, we introduce four auxiliary matrices $A$, $C1$, $C2$, and $C3$, each of which is a three-dimensional matrix of size $v\times n_{d}\times n_{d}$. These four auxiliary matrices encompass $v$ perspectives (distinct feature sets within multi-view learning), with each perspective having matrices of identical row and column dimensions. Specifically, $A$ denotes the relationship matrix among the various views, $C1$ is the auxiliary matrix for different views after singular value thresholding, $C2$ is the auxiliary matrix for different views following sparsity processing, and $C3$ is the auxiliary matrix that integrates information from multiple views. These auxiliary matrices are initialized to zero. Concurrently, we incorporate four Lagrange multiplier matrices $\lambda 1\in\mathbb{R}^{(v+1)n_{d}\times n_{d}}$, $\lambda 2\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, $\lambda 3\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and $\lambda 4\in\mathbb{R}^{v\times n_{d}\times n_{d}}$, and set five hyperparameters $\mu$, $\beta_{1}$, $\beta_{2}$, $\lambda$, and $\mu_{\text{max}}$, with only the parameter $\mu$ being continuously updated. For all formulas in this section, the subscript $i$ signifies the index of a particular view, where $i=0,1,\ldots,v$. We define $C_{\text{sum}}$, which represents the sum of the values of the $C2$ matrices of all views except view $i$ itself.
$C_{\text{sum}_{i}}=\sum_{i^{\prime}\neq i}\text{C2}_{i^{\prime}}$ (16)
Then, we update the auxiliary matrix ${A}$ iteratively. It updates the main relationship matrix ${A}$ of each view by integrating the auxiliary matrices of all views and the Lagrange multiplier matrices. The formula is as follows:
${A}_{i}=\;$ $\mu_{i}(K+3I)^{-1}\cdot\Big{(}\mu_{i}(K+C1+C2+C3)$
$-(\lambda 2+\lambda 3+\lambda 4)+Y_{i}^{T}\lambda 1\Big{)}$ (17)
where $K=Y_{i}^{T}Y_{i}$, and $I$ is the identity matrix. Subsequently, we update the auxiliary matrix $C1$ iteratively and perform singular value thresholding on it. This process can effectively reduce the dimensionality and denoise the matrix. We perform singular value decomposition:
${A}_{i}+\frac{\lambda 3}{\mu_{i}}=U_{i}\Sigma_{i}V_{i}^{T}$ (18)
where $U_{i}$ and $V_{i}$ are orthogonal matrices, and $\Sigma_{i}$ is the diagonal matrix containing the singular values. Then, the singular value matrix $\Sigma_{i}$ is thresholded to obtain $\Sigma_{i}^{\prime}=\max(\Sigma_{i}-\frac{\beta_{1}}{\mu_{i}},0)$. Using the thresholded singular value matrix to reconstruct the matrix $C1$, we get $C1_{i}=U_{i}\Sigma_{i}^{\prime}V_{i}^{T}$, and this completes the iterative update for $C1$.Next, we update $C2$ through soft thresholding, with the formula as follows:
$C2=\max({A}+\frac{\lambda 2}{\mu_{i}}-\frac{\beta_{2}}{\mu_{i}},0)+\min({A}+\frac{\lambda 2}{\mu_{i}}+\frac{\beta_{2}}{\mu_{i}},0)$ (19)
This operation updates the auxiliary matrix $C2$ for each view, effectively making the matrix sparser. Then, by integrating information from other matrices, we update the current matrix $C3$ to promote consistency and convergence, with the formula as follows:
$C3_{i}=\frac{2\lambda(v-1)C_{\text{sum}_{i}}+\mu_{i}{A}_{i}+\lambda 4_{i}}{2\lambda(v-1)+\mu_{i}}$ (20)
Subsequently, we adjust the Lagrange multipliers incrementally through the differences in constraint conditions, ensuring that the final solution satisfies all constraint conditions. The formula for this is as follows:
[…]
In the iterative optimization algorithm, convergence checking is an important step to determine whether the algorithm has achieved the expected precision, thus allowing the iteration to be terminated. Specifically, the goal of convergence checking is to ensure that the differences (errors) between various matrices are within an allowable error threshold. In the model, convergence checking is implemented by calculating the error values between different matrices, with the formulas as follows:
[…]
If all errors are less than the threshold value $\varepsilon=0.000001$, the algorithm is considered to have converged. It should be noted that during one iteration, $i$ starts from zero and gradually increases to the number of views $v$, with matrix updates performed at each step. That is, in one iteration, the matrices mentioned earlier are all updated $v$ times. When one iteration ends, $i$ is reset to zero and the process of incrementing and updating matrices begins again. The iteration stops after 100 iterations or if the algorithm does not converge. As previously mentioned, $\mu$ is also updated with each iteration. In the iterative optimization algorithm, updating $\mu$ serves to control the weight of different terms in the objective function, allowing for a gradual approach to the optimal solution with each iteration. Specifically, $\mu$ is a regularization parameter that is updated according to certain rules in each iteration. We set a maximum value $\mu_{\text{max}}$, and an increment factor $\rho$ to update the value of $\mu$:
$\mu_{t+1}=\min(\rho\mu_{t},\mu_{\text{max}})$ (23)
where $\rho$ is a constant greater than 1, and $\mu_{\text{max}}$ is the upper limit of $\mu$. Once all iterations are completed, we obtain the final auxiliary matrix $C2$ after the iterative updates. We then define a matrix $C_{\text{avg}}$, which represents the average of the $C2$ matrices from all views:
$C_{\text{avg}}=\frac{1}{v}\sum_{j}C2_{j}$ (24)
Ultimately, through pairwise multi-view learning on multiple drug features, we derive the drug affinity matrix $A_{DD}=|C_{\text{avg}}|+|C^{T}_{\text{avg}}|$. Similarly, we can obtain $A_{TT}$ using the aforementioned method. It is important to note that we use the iLearn software package to calculate features for the target sequences in these four different aspects (ACC, MORAN, CDT, PAAC), hence the number of views at this time is $v=v_{1}+v_{2}+v_{3}+v_{4}$. Subsequently, we apply the min-max normalization method to both $A_{DD}$ and $A_{TT}$.
Table 4: Performance comparison of SOC-DGL with baseline methods on four datasets.
Dataset Situation Model Year AUROC AUPR F1_score ACC Recall Specificity Precision
KIBA balanced GraphDTA 2021 0.9295 0.9355 0.8783 0.8566 0.9237 0.7710 0.8372
LRSpNM 2021 0.9224 0.9382 0.8740 0.8740 0.8740 0.7730 0.8371
MLLMC 2022 0.9484 0.9592 0.9049 0.8935 0.9038 0.8803 0.9059
MULGA 2023 0.9479 0.9551 0.8984 0.8773 0.9404 0.7911 0.8600
MMDG-DTI 2025 0.9586 0.9646 0.9083 0.8949 0.9318 0.8483 0.8858
SOC-DGL \ 0.9761 0.9805 0.9350 0.9268 0.9373 0.9133 0.9328
imbalanced GraphDTA 2021 0.9195 0.6500 0.6121 0.8911 0.6895 0.9198 0.5502
LRSpNM 2021 0.9132 0.6904 0.6339 0.8998 0.6900 0.9299 0.5863
MLMC 2022 0.9329 0.7462 0.6874 0.9192 0.7084 0.9494 0.6675
MULGA 2023 0.8918 0.5556 0.5628 0.8666 0.6808 0.8934 0.4796
MMDG-DTI 2025 0.9245 0.7193 0.6639 0.9102 0.6998 0.9391 0.6273
SOC-DGL \ 0.9364 0.7473 0.6908 0.9207 0.7103 0.9506 0.6730
Davis balanced GraphDTA 2021 0.9015 0.9324 0.8707 0.8330 0.9016 0.7195 0.8418
LRSpNM 2021 0.8930 0.9256 0.8644 0.8255 0.9083 0.6951 0.8250
MLMC 2022 0.9072 0.9395 0.8712 0.8712 0.9098 0.7177 0.8356
MULGA 2023 0.9158 0.9451 0.8790 0.8561 0.8539 0.8596 0.9058
MMDG-DTI 2025 0.9356 0.9647 0.8905 0.8627 0.9057 0.7934 0.8758
SOC-DGL \ 0.9467 0.9673 0.9038 0.8816 0.9097 0.8372 0.8985
imbalanced GraphDTA 2021 0.8486 0.6024 0.5514 0.8797 0.5096 0.9425 0.6006
LRSpNM 2021 0.9108 0.7067 0.6415 0.8961 0.6348 0.9409 0.6517
MLMC 2022 0.9110 0.7253 0.6704 0.9057 0.6530 0.9492 0.6888
MULGA 2023 0.8812 0.6179 0.5710 0.8698 0.5874 0.9187 0.5556
MMDG-DTI 2025 0.9087 0.7787 0.7369 0.9326 0.6503 0.9805 0.8409
SOC-DGL \ 0.9192 0.7655 0.6918 0.9134 0.6642 0.9561 0.7238
BindingDB balanced GraphDTA 2021 0.9325 0.9271 0.8719 0.8605 0.9103 0.8063 0.8366
LRSpNM 2021 0.8514 0.5771 0.5533 0.8706 0.5464 0.9263 0.5602
MLMC 2022 0.8797 0.9285 0.8707 0.8750 0.8237 0.9286 0.9233
MULGA 2023 0.9874 0.9859 0.9678 0.9660 0.9818 0.9490 0.9542
MMDG-DTI 2025 0.9620 0.9661 0.9181 0.9172 0.9172 0.9340 0.9340
SOC-DGL \ 0.9894 0.9888 0.9641 0.9627 0.9677 0.9574 0.9606
imbalanced GraphDTA 2021 0.9718 0.8099 0.7614 0.9320 0.7724 0.9718 0.7508
LRSpNM 2021 0.8481 0.5727 0.5866 0.9096 0.6383 0.9399 0.5426
MLMC 2022 0.8992 0.8260 0.7990 0.9617 0.7584 0.9844 0.8443
MULGA 2023 0.9706 0.8366 0.7188 0.9444 0.6991 0.9721 0.7395
MMDG-DTI 2025 0.9149 0.7428 0.6967 0.8980 0.6504 0.9508 0.7429
SOC-DGL \ 0.9790 0.8705 0.8057 0.9594 0.8406 0.9726 0.7744
DrugBank balanced GraphDTA 2021 0.8762 0.8676 0.8304 0.8033 0.9242 0.6719 0.7539
LRSpNM 2021 0.7563 0.8206 0.7016 0.6687 0.7522 0.5790 0.6573
MLMC 2022 0.8637 0.9126 0.8371 0.8415 0.7828 0.9052 0.8995
MULGA 2023 0.9715 0.9743 0.9165 0.9141 0.9112 0.9171 0.9219
MMDG-DTI 2025 0.9626 0.9658 0.9180 0.9159 0.9083 0.9240 0.9281
SOC-DGL \ 0.9735 0.9774 0.9286 0.9267 0.9208 0.9331 0.9369
imbalanced GraphDTA 2021 0.8450 0.4268 0.4564 0.8764 0.5336 0.9133 0.3987
LRSpNM 2021 0.7752 0.5480 0.5437 0.9257 0.4577 0.9758 0.6695
MLMC 2022 0.8594 0.7330 0.7043 0.9498 0.6181 0.9853 0.8185
MULGA 2023 0.9259 0.7179 0.6667 0.9358 0.6579 0.9659 0.6756
MMDG-DTI 2025 0.9051 0.6978 0.6717 0.9381 0.6522 0.9688 0.6943
SOC-DGL \ 0.9295 0.7654 0.7211 0.9508 0.6567 0.9824 0.8004
6.3 Comparison with Baseline Methods in “warm-start” experiments
To validate the effectiveness and superiority of SOC-DGL, we conducted comparative experiments with four leading methods that were widely recognized in DTI prediction. These models included GraphDTA [27], LRSpNM [28], MLMC [29], MULGA [30], and MMDG-DTI [31]. The detailed results of the comparative experiments on these baseline datasets were presented in Appendix Tab. 4.
6.3.1 GraphDTA
GraphDTA, proposed by Thin Nguyen et al. in 2021, is a graph neural network-based model for predicting drug-target binding affinity. It represents drug molecules as graphs, using graph convolution to capture their spatial and chemical properties, improving DTI prediction accuracy by integrating various drug and target features.
6.3.2 LRSpNM
LRSpNM, introduced by Gaoyan Wu et al. in 2021, predicts drug-target interactions using similarity information and dynamic pre-filling of unknown interactions. It employs low-rank matrix approximation, loss functions, and Laplacian regularization, utilizing the ADMM algorithm for efficient computation on large bioinformatics datasets.
6.3.3 MLMC
MLMC, proposed by Yixin Yan et al. in 2022, predicts drug-disease associations by integrating multi-view learning for similarity information and matrix completion to fill missing values in the association matrix, improving the model’s predictive power for unknown drug-disease pairs.
6.3.4 MULGA
MULGA, proposed by Jiani Ma et al. in 2023, is a multi-view graph autoencoder framework for DTI and drug repurposing prediction. It learns affinity matrices and infers missing interactions using a novel “guilt-by-association” negative sampling method, improving predictive accuracy.
6.3.5 MMDG-DTI
MMDG-DTI, proposed by Yang Hua et al. in 2025, addresses domain-specific biases in existing models by using pre-trained LLMs to extract generalized features and a hybrid GNN for structural learning. It enhances generalization through Domain Adversarial Training, contrastive learning, and multi-modal feature fusion.
6.4 Analysis of the even-polynomial graph filter in EDGL
Higher-order homogeneous information enhancement by the even-polynomial graph filter was primarily designed to capture direct relationships between adjacent nodes. In the context of a drug-target network, this approach emphasized identifying the direct interactions between drugs and targets. It effectively highlighted the interactions between first-order neighbors, which were nodes directly connected by an edge. Higher-order homogeneous information enhancement by the even-polynomial graph filter retained odd-order monomials while discarding even-order terms, as formulated below:
$Y^{\prime}=f(\hat{X},P)=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k-1}\hat{X}$ (25)
Regarding the self-attention mechanism, its application in graph neural networks captured the complex interactions between nodes in a graph. By calculating attention weights, the self-attention mechanism assigned different weights to different neighboring nodes, thereby selectively aggregating important neighboring information. It focused on the most critical relationships for the prediction task, whether direct or indirect. For the self-attention mechanism, we used the result $H^{\prime}$ obtained from ADGL as the input and performed three different linear transformations on the input vector to obtain Q (Query), K (Key), and V (Value), where $Q=H^{\prime}W^{Q},K=H^{\prime}W^{K},V=H^{\prime}W^{V}$. Here, $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times d_{k}}$ were trainable weight matrices, and here $d_{k}=m$. Then, the formula for self-attention was obtained as:
$Y^{\prime\prime}=\text{softmax}(\frac{QK^{T}}{\sqrt{m}})\cdot V$ (26)
6.5 Validation experiment on imbalanced loss function in imbalanced datasets
The standard binary cross-entropy loss function was a commonly used loss function suitable for balanced datasets, but it might have caused the model to favor the majority class in imbalanced datasets. The focal loss [51] aimed to address class imbalance by reducing the weight of easily classified samples while increasing the weight of hard-to-classify samples. The weighted binary cross-entropy loss function balanced the disparity between positive and negative samples by assigning higher weights to the minority class. The reduced negative weighted binary cross-entropy loss function reduced the weight of negative samples while keeping the weight of positive samples unchanged, mitigating the impact of imbalanced data on model performance. The formulas for the four loss functions were as follows:
$\mathcal{L}_{S}=-\frac{1}{n_{d}\cdot n_{t}}(\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (27)
$\mathcal{L}_{F}=-\frac{|\mathcal{Y}^{-}|}{|\mathcal{Y}^{+}|}\left(1-h_{ij}^{*}\right)^{\gamma}\log(h_{ij}^{*})$ (28)
$\mathcal{L}_{W}=-\frac{1}{n_{d}\cdot n_{t}}(\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (29)
$\mathcal{L}_{R}=-\frac{1}{n_{d}\cdot n_{t}}(\varpi\frac{\mid y^{-}\mid}{\mid y^{+}\mid}\!\!\sum_{(i,j)\in y^{+}}\!\!\!\!\log(h_{ij}^{*})+\!\!\!\!\sum_{(i,j)\in y^{-}}\!\!\!\!\log(1-h_{ij}^{*}))$ (30)
where $\mathcal{L}_{S}$, $\mathcal{L}_{F}$, $\mathcal{L}_{W}$, and $\mathcal{L}_{R}$ were Standard Binary Cross-Entropy Loss (SLF), Focal Loss with Dynamic Scaling (FLF), Weighted Binary Cross-Entropy Loss (WLF), and Weighted Binary Cross-Entropy Loss with Reduced Negative Samples (RLF), respectively. $n_{d}$ was the number of drugs in the dataset, $n_{t}$ was the number of targets in the dataset, $|y^{+}|$ represented the number of positive samples, and $|y^{-}|$ represented the number of negative samples. $h_{ij}^{*}$ represented the model’s predicted probability of interaction between drug $i$ and target $j$. $\gamma$ was an adjustment parameter used to reduce the weight of easily classified samples and increase the weight of hard-to-classify samples. $\varpi$ was also an adjustment parameter, used to reduce the weight of negative samples while keeping the weight of positive samples unchanged.